Circular No. 6894

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 J1 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery by S. Stezelberger of a bright comet in the northwestern corner of the field of view of the C3 coronagraph, ∼ 8^{o} from the sun:

1998 UT  α_{2000}  δ_{2000}
May 3.779  2^{h}10.^{m}2  +19^{o}35′

Preliminary measurements by Stezelberger and St. Cyr, reduced by G. V. Williams and given on MPEC 1998-J13, show the comet, estimated by St. Cyr at mag 0 and possibly brighter, continuing to move across the top of the C3 field. There is a fanshaped tail extending antisunward at least to the edge of the field, and there may be a separate narrow tail extending roughly along the direction of motion. Computations by the undersigned show that the observations are represented rather unequivocally by the following parabolic orbital elements and suggest that the comet might become observable from the ground, particularly later in the month at solar elongations favoring the southern hemisphere.

T = 1998 May 8.901 TT  ω = 107.^{o}577
Ω = 348.442 }2000.0
q = 0.17454 AU  i = 55.085
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 5 2^{h}23.^{m}55 +21^{o}24.^{′}8 1.190 0.234 7.^{o}7 35.^{o}4  0.1
6 2 35.00 +22 50.6 1.165 0.211  7.5  38.6 −0.4
7 2 47.69 +24 06.1 1.137 0.191  7.6  44.3 −0.9
8 3 01.57 +25 04.1 1.107 0.178  8.1  52.9 −1.3
9 3 16.29 +25 37.0 1.075 0.175  8.9  63.6 −1.4
10 3 31.27 +25 40.0 1.042 0.180  9.9  74.8 −1.4
11 3 45.90 +25 13.5 1.010 0.194 11.0  84.5 −1.1
12 3 59.81 +24 22.3 0.980 0.215 12.3  91.9 −0.7
13 4 12.86 +23 12.2 0.953 0.239 13.6  96.9 −0.3
14 4 25.07 +21 48.3 0.930 0.266 15.0 100.0  0.1
15 4 36.54 +20 14.3 0.910 0.294 16.6 101.6  0.5
16 4 47.35 +18 33.2 0.892 0.323 18.2 102.2  0.8
17 4 57.60 +16 47.0 0.878 0.352 19.9 101.9  1.2

1998 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden